OpenStudy (anonymous):

6. [6.02] Johnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school? (4 points) 22 42 51 60 7. [6.02] Solve the following system. y = x + 3 3x + y = 19 (4 points) (7, 4) (–4, 7) (4, 7) (4, –7) 8. [6.03] What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −1 (4 points) –4 4 –3 3

OpenStudy (anonymous):

will u skype me if i help u girl

OpenStudy (anonymous):

Yeaaah sure!

OpenStudy (anonymous):

@Yashhere

OpenStudy (anonymous):

im doing it babe

OpenStudy (anonymous):

Kay thanks.

OpenStudy (anonymous):

I have more for you If you want! PM me

OpenStudy (anonymous):

listen up girl 22 (4,7) 4

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

9. [6.03] Solve the following system. x − y = 10 x + y = 6 (5 points) (–2, 8) (–8, 2) (8, –2) (8, 2) 10. [6.03] If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 3x + y = 6 4x − 5y = 10 (4 points) Multiply the first equation by 5. Multiply the first equation by –5. Multiply the first equation by –4. Add the two equations together. 11. [6.03] Auditions for a new, upcoming band in Chicago were held with 72 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 34, how many women auditioned? (5 points) 22 44 50 60 12. [6.05] Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $21.95 per month plus an additional $0.07 per minute. Cell Best has a plan that costs $25.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month? (5 points) 0.12m = 4.00 21.95 + 0.07m = 26.00 21.95 + 0.07m = 25.95 + 0.05m 25.95 + 0.05m = 22.02

OpenStudy (anonymous):

are u pelletting me

OpenStudy (anonymous):

Noo

OpenStudy (anonymous):

skype me or u will be killed

OpenStudy (anonymous):

After you answer those.

OpenStudy (anonymous):

we already had a deal

OpenStudy (anonymous):

mmkay Answer them.

OpenStudy (anonymous):

u skyping

OpenStudy (kamille):

hahahha

OpenStudy (anonymous):

Hey can you help? @Kamille

OpenStudy (anonymous):

i helped her but what did she do blocked me i hate such people so much

OpenStudy (anonymous):

I just need the other questions answered too. @Yashhere

OpenStudy (anonymous):

but uve already blocked me

OpenStudy (anonymous):

I unblocked you maan.

OpenStudy (kamille):

x − y = 10 x + y = 6 As you can see, y can be eliminitaded by adding one eq to other.So you get 2x=16 x=8 now put x in other eq and you will get Y 8-y=10 -y=2 y=-2 answer is (8;-2)

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

she did the easy one

OpenStudy (anonymous):

I need them ALL.

OpenStudy (kamille):

second one is Multiply the first equation by 5.

OpenStudy (anonymous):

Alright thanks..

OpenStudy (anonymous):

More?

OpenStudy (anonymous):

arent u done skype me

OpenStudy (anonymous):

I have like 12 more questions.

OpenStudy (anonymous):

screw them skype is more appealing

OpenStudy (anonymous):

I dont care, I need this done. Its 13 days late and I have to pass it,

OpenStudy (anonymous):

Otherwise NO COLLEGE FOR ME!

OpenStudy (anonymous):

which grade u in

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

Just help OH MY GOD.

OpenStudy (anonymous):

and u are solving a 5 grade question

OpenStudy (anonymous):

Im in standard algebra, yep

OpenStudy (anonymous):

This is all review, of extra credit.

OpenStudy (anonymous):

u skype i tell u there

OpenStudy (anonymous):

ITs @vividaisy. If you dont I willl block you

OpenStudy (anonymous):

are u ashley

OpenStudy (anonymous):

mhhm

OpenStudy (anonymous):

accept it

OpenStudy (anonymous):

u are such a cheater

OpenStudy (anonymous):

@ThatHippieKid will u skype me!!!!!!!LOL jk